# TF-Mamba: Text-enhanced Fusion Mamba with Missing Modalities for Robust Multimodal Sentiment Analysis

### v1

@@ Abstract
Multimodal Sentiment Analysis (MSA) with missing modalities has attracted increasing attention recently. While current Transformer-based methods leverage dense text information to maintain model robustness, their quadratic complexity hinders efficient long-range modeling and multimodal fusion. To this end, we propose a novel and efficient Text-enhanced Fusion Mamba (TF-Mamba) framework for robust MSA with missing modalities. Specifically, a Text-aware Modality Enhancement (TME) module aligns and enriches non-text modalities, while reconstructing the missing text semantics. Moreover, we develop Text-based Context Mamba (TC-Mamba) to capture intra-modal contextual dependencies under text collaboration. Finally, Text-guided Query Mamba (TQ-Mamba) queries text-guided multimodal information and learns joint representations for sentiment prediction. Extensive experiments on three MSA datasets demonstrate the effectiveness and efficiency of the proposed method under missing modality scenarios. Our code is available at https://github.com/codemous/TF-Mamba.
@@
Xiang Li1, Xianfu Cheng2, Dezhuang Miao1, Xiaoming Zhang1, Zhoujun Li2 1School of Cyber Science and Technology, Beihang University, Beijing 2School of Computer Science and Engineering, Beihang University, Beijing 1,2{xlggg, buaaxcf, taiyue, yolixs, lizj}@buaa.edu.cn
@@ 1 Introduction
Recent studies (Zhao et al., 2021; Yuan et al., 2021, 2023; Li et al., 2024b; Zhang et al., 2024) attempt to tackle the challenge of missing modalities in MSA. Among them, Transformer-based fusion models achieve notable progress owing to their powerful sequence modeling capabilities. For example, TFR-Net (Yuan et al., 2021) adopts a Transformer-based feature reconstruction strategy to recover missing information in multimodal sequences. More recently, LNLN (Zhang et al., 2024) prioritizes high-quality text features and treats it as the dominant modality to improve model robustness under various noise conditions. Unfortunately, these approaches often suffer from quadratic computational complexity and fail to achieve efficient text-enhanced multimodal fusion. Mamba (Gu and Dao, 2023) emerges as a linear-time alternative to Transformers, exhibiting great promise in various domains (Zhu et al., 2024; Yang et al., 2024; Li et al., 2024d). Nevertheless, incorporating text enhancement into Mamba-based architectures for efficient missing modality modeling and fusion remains unexplored, which is crucial for enhancing both the performance and robustness of MSA.
@@
Extensive experiments on three MSA benchmarks demonstrate the superiority of TF-Mamba under uncertain missing modality scenarios. For instance, on MOSI, TF-Mamba outperforms Transformer-based state-of-the-art with 90% fewer parameters and FLOPs.
@@ 2.1 Multimodal Sentiment Analysis
Multimodal Sentiment Analysis (MSA) integrates multimodal information to understand and analyze human sentiments. Mainstream studies in MSA (Wu et al., 2021; Yang et al., 2023; Li et al., 2025; Sun and Tian, 2025) focus on developing sophisticated fusion strategies and interaction mechanisms to improve sentiment prediction performance. For example, Han et al. (2021) hierarchically maximize mutual information between unimodal pairs to enhance multimodal fusion. Zhang et al. (2023) leverage sentiment-intensive cues to guide the representation learning of other modalities. Despite these progresses, most methods assume fully available modalities, which is rarely achievable in real-world scenarios due to random data missing issues.
@@ 2.2 State Space Models and Mamba
Building on Mamba’s success, several studies (Eom et al., 2024; Li et al., 2024d, a; Qiao et al., 2024; Li et al., 2024e) explore its potential for multimodal fusion. For instance, Dong et al. (2024) integrate features from different modalities into unimodal Mamba networks for cross-modal interactions. Qiao et al. (2024) concatenate visual and textual sequences before jointly modeling them through Mamba blocks. Li et al. (2024e) learns both local and global cross-modal alignments prior to Mamba-based multimodal fusion. Recently, Ye et al. (2025) pioneer the use of Mamba for depression detection, combining hierarchical modeling with progressive fusion strategies. Overall, these Mamba-based frameworks deliver competitive performance and improved efficiency compared to Transformer-based methods. However, its application to MSA with missing modalities remains underexplored, especially in integrating text-enhanced fusion strategies into Mamba’s efficient modeling architecture. In contrast, we treat text as dominant modality during Mamba-based sequence modeling and multimodal fusion to enhance model robustness in incomplete modality settings.
@@ 3.4 Text-aware Modality Enhancement
We first unify sequence length and feature dimension using a CTC-based temporal alignment (Tsai et al., 2019; Zhou et al., 2024) with an MLP:
@@
The same way is applied to enhance the audio modality, resulting in enriched audio representations . Given the rich sentiment semantics in text, we reconstruct missing textual information to enhance model robustness. A simple MLP-based decoder , consisting of two linear layers and a ReLU activation, is adopted as the reconstructor. We employ text Smooth L1 loss (Yuan et al., 2021; Sun et al., 2023) to evaluate the reconstruction:
@@ 3.5 Text-based Context Mamba
where , , , , and , represent the input, hidden state, and output features at time step. The backward SSMs follows an analogous structure using the , , and matrixs. The overall process of Bi-Mamba is formally as:
@@ 4.5 Ablation Study
When the text enhancement strategy is removed, model performance degrades notably, which underscores the effectiveness of text-enhanced fusion for sentiment prediction. Similarly, without the text reconstruction loss, the model struggles to recover missing textual semantics, leading to poorer multimodal representations. This highlights the importance of reconstructing incomplete text features for robust MSA. We further evaluate model performance under complete modality missing conditions (), with results and discussions summarized in Appendix D. The complete absence of text modality causes a marked performance drop, notably when paired with missing audio or visual inputs. In contrast, missing only audio or visual modality has a limited impact. These findings further underscore the significance of text enhancement in achieving robust MSA.
@@
Removing the TME module leads to consistent performance drops on both datasets. These results show that TME effectively facilitates representation learning and multimodal fusion through text-aware alignment and enhancement. In addition, incorporating the TME module into the baseline BI-Mamba improves ACC-7 by 3.26% on MOSI and ACC-5 by 2.13% on SIMS. Such outcomes substantiate both the effectiveness and rationality of the TME module.
@@ 4.6 Efficiency Study
We conduct an efficiency analysis of TF-Mamba, focusing on its intra-modal modeling and cross-modal interaction components. To ensure fair comparisons, we exclude the computational cost of pre-trained encoders. All experiments are performed under identical conditions, evaluating model parameters (Params), floating-point operations (FLOPs), and performance metrics (ACC-7 and F1) on the MOSI dataset. Comparisons are made against the Transformer-based SOTA baseline LNLN and a TF-Mamba variant, TF-Trans, where Mamba blocks are replaced with Transformer blocks.
@@ Appendix A Feature Extraction
To ensure fair and consistent comparisons with existing SOTA methods, we adopt the official released features provided by the corresponding benchmark datasets, following the unified MMSA framework (Mao et al., 2022). All features are pre-extracted for text, audio, and visual modalities using standard toolkits, as detailed below.
@@ Appendix C Details of Robust Comparison
Tables 9, 10, and 11 present detailed robustness comparisons on the MOSI, MOSEI, and SIMS datasets, respectively. We observe that when the missing rate is low, Self-MM achieves notable advantages across several evaluation metrics. However, as increases, TF-Mamba consistently outperforms other methods on most metrics, demonstrating its ability to learn robust multimodal representations under varying levels of data incompleteness. Additionally, while models such as LNLN and TF-Mamba perform well under high missing modality rates, they often struggle to maintain optimal performance when modality missingness is low. Balancing robustness and accuracy across different missing conditions remains challenging.
@@ Appendix D Analysis of Complete Modality Missing
We conduct comprehensive experiments under complete modality missing conditions to further assess model robustness and examine the impact of discarding different modalities. The detailed results on the MOSI, MOSEI, and SIMS datasets are presented in Tables 6, 7, and 8, respectively.
@@
| BI-Mamba | 38.78 | 42.71 | 81.25 / 80.03 | 81.34 / 80.06 | 0.873 | 0.699 |
| LNLN | 39.26 | 44.70 | 80.99 / 78.96 | 80.86 / 78.92 | 0.834 | 0.714 |
| TF-Mamba | 42.86 | 48.40 | 81.86 / 80.03 | 81.87 / 79.97 | 0.824 | 0.732 |
| Random Missing Rate | ||||||
| MISA | 34.60 | 38.97 | 75.76 / 74.54 | 75.82 / 74.51 | 0.989 | 0.618 |
| Self-MM | 34.89 | 40.67 | 76.37 / 74.98 | 75.68 / 74.94 | 0.967 | 0.614 |
| MMIM | 36.83 | 40.43 | 74.08 / 71.91 | 73.47 / 71.28 | 0.974 | 0.612 |
| TFR-Net | 32.55 | 38.34 | 72.36 / 71.28 | 70.12 / 69.58 | 1.065 | 0.572 |
| CENET | 34.74 | 38.97 | 76.83 / 72.01 | 76.56 / 71.30 | 0.983 | 0.605 |
| ALMT | 33.04 | 37.17 | 75.15 / 72.94 | 75.51 / 73.66 | 0.992 | 0.596 |

### v2

@@ Abstract
Multimodal Sentiment Analysis (MSA) with missing modalities has attracted increasing attention recently. While current Transformer-based methods leverage dense text information to maintain model robustness, their quadratic complexity hinders efficient long-range modeling and multimodal fusion. To this end, we propose a novel and efficient Text-enhanced Fusion Mamba (TF-Mamba) framework for robust MSA with missing modalities. Specifically, a Text-aware Modality Enhancement (TME) module aligns and enriches non-text modalities, while reconstructing the missing text semantics. Moreover, we develop Text-based Context Mamba (TC-Mamba) to capture intra-modal contextual dependencies under text collaboration. Finally, Text-guided Query Mamba (TQ-Mamba) queries text-guided multimodal information and learns joint representations for sentiment prediction. Extensive experiments on three MSA datasets demonstrate the effectiveness and efficiency of the proposed method under missing modality scenarios. Code is available at https://github.com/codemous/TF-Mamba.
@@
Xiang Li1††thanks: Equal contribution., Xianfu Cheng211footnotemark: 1, Dezhuang Miao1, Xiaoming Zhang1††thanks: Corresponding author., Zhoujun Li2 1School of Cyber Science and Technology, Beihang University, Beijing 2School of Computer Science and Engineering, Beihang University, Beijing {xlggg, buaacxf, taiyue, yolixs, lizj}@buaa.edu.cn
@@ 1 Introduction
Recent studies (Yuan et al., 2021, 2023; Li et al., 2024b; Zhang et al., 2024) attempt to tackle the challenge of missing modalities in MSA. Among them, Transformer-based fusion models achieve notable progress owing to their powerful sequence modeling capabilities. For example, TFR-Net (Yuan et al., 2021) adopts a Transformer-based feature reconstruction strategy to recover missing information in multimodal sequences. More recently, LNLN (Zhang et al., 2024) prioritizes high-quality text features and treats text as the dominant modality to improve model robustness under various noise conditions. Unfortunately, these approaches often suffer from quadratic computational complexity and fail to achieve efficient text-enhanced multimodal fusion. Mamba (Gu and Dao, 2023) emerges as a linear-time alternative to Transformers, exhibiting great promise in various domains (Yang et al., 2024; Li et al., 2024d). Nevertheless, incorporating text enhancement into Mamba-based architectures for efficient missing modality modeling and fusion remains unexplored, which could potentially contribute to enhancing MSA performance.
@@
Extensive experiments on three MSA benchmarks demonstrate the superiority of TF-Mamba under uncertain missing modality scenarios. For instance, on MOSI, TF-Mamba outperforms Transformer-based state-of-the-art with fewer parameters and FLOPs.
@@ 2.1 Multimodal Sentiment Analysis
Multimodal Sentiment Analysis (MSA) integrates multimodal information to understand and analyze human sentiments. Mainstream studies in MSA (Li et al., 2024e, 2025a, 2025b; Sun and Tian, 2025) focus on developing sophisticated fusion strategies and interaction mechanisms to improve sentiment prediction performance. For example, Han et al. (2021) hierarchically maximize mutual information between unimodal pairs to enhance multimodal fusion. Zhang et al. (2023) leverage sentiment-intensive cues to guide the representation learning of other modalities. Despite these progresses, most methods assume fully available modalities, which is rarely achievable in real-world scenarios due to random data missing issues.
@@ 2.2 State Space Models and Mamba
Building on Mamba’s success, several studies (Li et al., 2024d, a; Dong et al., 2025; Li et al., 2025c) explore its potential for multimodal fusion. For instance, Qiao et al. (2024) concatenate visual and textual sequences before jointly modeling them through Mamba blocks. Li et al. (2025c) learns both local and global cross-modal alignments prior to Mamba-based multimodal fusion. Recently, Ye et al. (2025) pioneer the use of Mamba for depression detection, combining hierarchical modeling with progressive fusion strategies. Overall, these Mamba-based frameworks deliver competitive performance and improved efficiency compared to Transformer-based methods. However, its application to MSA with missing modalities remains underexplored, especially in integrating text-enhanced fusion strategies into Mamba’s efficient modeling architecture. In contrast, we treat text as dominant modality during Mamba-based sequence modeling and multimodal fusion to enhance model robustness in incomplete modality settings.
@@ 3.4 Text-aware Modality Enhancement
We first unify sequence length and feature dimension using a CTC-based temporal alignment (Zhou et al., 2024) with a linear mapping:
@@
The same way is applied to enhance the audio modality, resulting in enriched audio representations . Given the rich sentiment semantics in text, we reconstruct missing textual information to enhance model robustness. A simple MLP-based decoder , consisting of two linear layers and a ReLU activation, is adopted as the reconstructor. We employ Smooth L1 loss (Yuan et al., 2021; Sun et al., 2023) to evaluate the text reconstruction:
@@ 3.5 Text-based Context Mamba
where , , , , and , represent the input, hidden state, and output features at time step. The backward SSMs follows an analogous structure using the , , and matrices. The overall process of Bi-Mamba is formally as:
@@ 4.5 Ablation Study
When the text enhancement strategy is removed, model performance degrades notably, which underscores the effectiveness of text-enhanced fusion for sentiment prediction. Similarly, without the text reconstruction loss, the model struggles to recover missing textual semantics, leading to poorer multimodal representations. This highlights the importance of reconstructing incomplete text features for robust MSA. We further evaluate model performance under complete modality missing conditions (), with results and discussions summarized in Appendix D. The complete absence of the text modality causes a marked performance drop, even when combined with missing audio or visual inputs. In contrast, missing only audio or visual modality has a limited impact. These findings further underscore the significance of text enhancement in robust MSA.
@@
Removing the TME module leads to consistent performance drops on both datasets. These results show that TME effectively facilitates representation learning and multimodal fusion through text-aware alignment and enhancement. In addition, incorporating the TME module into the baseline BI-Mamba improves Acc-7 by 4.23% on MOSI and Acc-5 by 2.13% on SIMS. Such outcomes substantiate both the effectiveness and rationality of the TME module.
@@ 4.6 Efficiency Study
We conduct an efficiency analysis of TF-Mamba, focusing on its intra-modal modeling and cross-modal interaction components. To ensure fair comparisons, we exclude the computational cost of pre-trained encoders. All experiments are performed under identical conditions, evaluating model parameters (Params), floating-point operations (FLOPs), and performance (Acc-7 and F1) on MOSI dataset. Comparisons are made against the Transformer-based SOTA baseline LNLN and a TF-Mamba variant, TF-Trans, where Mamba blocks are replaced with Transformer blocks.
@@
| TC-Mamba | TQ-Mamba | MAE | F1 | Acc-5 |
| 1 | 1 | 0.521 | 71.18 | 33.83 |
| 1 | 2 | 0.512 | 72.20 | 34.46 |
| 1 | 3 | 0.523 | 69.36 | 34.31 |
| 2 | 1 | 0.522 | 70.46 | 33.04 |
| 2 | 2 | 0.508 | 70.97 | 34.09 |
| 2 | 3 | 0.518 | 70.15 | 33.85 |
4.7 Sensitivity Analysis
In this section, we analyze the sensitivity of key hyperparameters on model performance. We focus on two factors: (1) the reconstruction loss weight, controlled by , and (2) the number of layers in the TC-Mamba and TQ-Mamba modules, which govern model capacity and computational cost.
Effect of Loss Weight
We first examine the influence of the reconstruction loss weight on model performance. This hyperparameter controls the relative importance of the reconstruction loss in guiding the model to recover missing semantic information. We evaluated the model across on the MOSI dataset, and the results are visualized in Figure 5. The analysis reveals a clear trend: excessively small values (e.g., 0.1) provide insufficient supervision for reconstructing missing semantics, leading to suboptimal sentiment prediction. Conversely, excessively large values (e.g., 0.9) cause the reconstruction loss to dominate, which interferes with the model’s primary sentiment inference. A gradual increase in leads to consistent gains in F1 and Acc-7, with the overall best performance observed at . This choice strikes an effective balance between reconstruction and sentiment prediction, underscoring the necessity of tuning appropriately.
Effect of Mamba Layers
We further investigate the impact of varying the number of TC-Mamba and TQ-Mamba layers on model performance. The evaluation on the SIMS dataset is summarized in Table 6. The results show that using too few layers fails to capture the complex cross-modal interactions, while excessive layers introduce higher computational cost without consistent performance gains and may even risk overfitting. Overall, performance improves as the number of layers increases, with the (1,2) configuration of TC-Mamba and TQ-Mamba achieving the best trade-off between efficiency and Accuracy, delivering strong results at a moderate computational cost.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (Nos. 62272025 and U22B2021), and by the Fund of the State Key Laboratory of Software Development Environment.
@@ Appendix A Feature Extraction
To ensure fair and consistent comparisons with existing SOTA methods, we adopt the officially released features provided by the corresponding benchmark datasets, following the unified MMSA framework (Mao et al., 2022). All features are pre-extracted for text, audio, and visual modalities using standard toolkits, as detailed below.
@@ Appendix C Details of Robust Comparison
Tables 10, 11, and 12 present detailed robustness comparisons on the MOSI, MOSEI, and SIMS datasets, respectively. We observe that when the missing rate is low, Self-MM achieves notable advantages across several evaluation metrics. However, as increases, TF-Mamba consistently outperforms other methods on most metrics, demonstrating its ability to learn robust multimodal representations under varying levels of data incompleteness. Additionally, while models such as LNLN and TF-Mamba perform well under high missing modality rates, they often struggle to maintain optimal performance when modality missingness is low. Balancing robustness and Accuracy across different missing conditions remains challenging.
@@ Appendix D Analysis of Complete Modality Missing
We conduct comprehensive experiments under complete modality missing conditions to further assess model robustness and examine the impact of discarding different modalities. The detailed results on the MOSI, MOSEI, and SIMS datasets are presented in Tables 7, 8, and 9, respectively.
@@
Appendix E Analysis of Threshold Selection in TME
In the TME module, a threshold is introduced to generate binary masks over the similarity matrix between text tokens and other modalities. This mechanism suppresses irrelevant information by filtering out weakly correlated token pairs. The threshold is set as the average similarity value , where denotes the number of text tokens. Token pairs with similarity scores below the threshold are masked, whereas those exceeding the threshold are retained, thereby mitigating the influence of weak correlations while ensuring the preservation of informative semantic relationships.
To assess robustness, we conduct a sensitivity analysis by varying the threshold within on the MOSI and SIMS datasets. The corresponding model performance is shown in Figure 7 and Figure 8. From Figure 7 (MOSI), a threshold of provides the overall best performance across F1 and Acc-7, whereas retains excessive irrelevant information and higher thresholds gradually remove important cues. Figure 8 (SIMS) exhibits a similar trend: consistently yields stable and optimal performance across both F1 and Acc-5, while degrades performance and overly high thresholds suppress too much semantic information. Overall, this analysis confirms as an effective threshold choice, striking a balance between noise suppression and semantic retention, and demonstrating consistent advantages across both datasets.
@@ Appendix E Analysis of Threshold Selection in TME
| MISA | 40.28 | 46.21 | 80.18 / 79.01 | 80.21 / 78.97 | 0.847 | 0.721 |
| Self-MM | 40.33 | 49.03 | 81.40 / 80.03 | 81.19 / 80.03 | 0.812 | 0.728 |
| MMIM | 42.61 | 46.65 | 79.98 / 78.13 | 79.83 / 77.99 | 0.825 | 0.718 |
| TFR-Net | 38.63 | 45.82 | 79.27 / 77.99 | 78.70 / 77.61 | 0.872 | 0.705 |
| CENET | 40.13 | 46.60 | 80.08 / 78.38 | 79.91 / 78.20 | 0.837 | 0.719 |
| ALMT | 39.84 | 45.48 | 80.90 / 78.67 | 81.15 / 79.08 | 0.843 | 0.703 |
| BI-Mamba | 38.78 | 42.71 | 81.25 / 80.03 | 81.34 / 80.06 | 0.873 | 0.699 |
| LNLN | 39.26 | 44.70 | 80.99 / 78.96 | 80.86 / 78.92 | 0.834 | 0.714 |
| TF-Mamba | 42.86 | 48.40 | 81.86 / 80.03 | 81.87 / 79.97 | 0.824 | 0.732 |
| Random Missing Rate | ||||||